OpenStudy (anonymous):

how do you solve: x squared - (x-2)(x-4)

OpenStudy (anonymous):

you can't solve this since this isn't an equation... best you can do with this expression is to simplify.

OpenStudy (mimi_x3):

\[x^2-(x-2)(x-4)\]??

OpenStudy (anonymous):

just simplify

OpenStudy (mimi_x3):

Well, first expand \((x-2)(x-4)\)

OpenStudy (anonymous):

yeah then what?

OpenStudy (mimi_x3):

After that collect the like terms. First what did you get when you expanded it?

OpenStudy (lgbasallote):

all you need shall be revealed once it has been expanded

OpenStudy (anonymous):

(x-2)(x-4) =x^2 -4x-2x +8 <---expand =x^2 -6x+8 x^2 - (x-2)(x-4) =x^2 - (x^2 -6x+8) =x^2-x^2+6x-8 =6x-8

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

understand now